OpenStudy (anonymous):

Prove that (1-w-w^2)^6=64

OpenStudy (anonymous):

U mean solve?

OpenStudy (anonymous):

yup

hartnn (hartnn):

i thin u can find what is w+w^2=??

OpenStudy (anonymous):

(+-2)^6=64

OpenStudy (anonymous):

Or u can use that

OpenStudy (anonymous):

how???

hartnn (hartnn):

1+w+w^2=0 what is w+w^2 ??

OpenStudy (anonymous):

where is w+w^2 applying

OpenStudy (anonymous):

(1-w-w^2)^6=64 (1-w-w^2)^6=(-+2)^6 1-w-w^2=-+2

OpenStudy (anonymous):

Now, can u do it from here?

hartnn (hartnn):

i think its given that w is the cube root of unity, isn't it ?

OpenStudy (anonymous):

(1-w-w^2)^6 should come 64

OpenStudy (anonymous):

not solving we have to prove

hartnn (hartnn):

1+w+w^2=0 so w+w^2= -1 LEFT SIDE = (1-(w+w^2))^6 =(1-(-1))^6 =2^6 =64 =RIGHT SIDE

OpenStudy (anonymous):

yeah that was i talking about